OpenStudy (anonymous):

2x^2 + 3x + 20 over x^2 - 5x - 36 my answer: 2(13x+5) over 5x -9

OpenStudy (anonymous):

pleassseeeee answer:(((((

OpenStudy (anonymous):

i need to study.

sam (.sam.):

Hold on

OpenStudy (anonymous):

ok:)

sam (.sam.):

Did you divide the top one with the bottom one?

sam (.sam.):

You will get something like this after synthetic division, \[\Large 2+\frac{13x+92}{x^2-5x-36}\] Then factor it, \[\Large \frac{13 x+92}{(x-9) (x+4)}\]

sam (.sam.):

We'll leave the "2" behind first, because we have to deal with partial fraction for that fraction. After partial fraction you should get this \[\frac{209}{13 (x-9)}-\frac{40}{13 (x+4)}\] Add back the "2", \[2+\frac{209}{13 (x-9)}-\frac{40}{13 (x+4)}\]

sam (.sam.):

Actually your answer doesn't equal to your question, you can check that by substituting lets say x=5 into it, For your question, if x=5, then y=-85/36 But for your answer, if x=5, then y=35/4 The 2 y's doesn't equal means your answer is wrong, this is just a method to check.

OpenStudy (anonymous):

thats not an option dear:(

sam (.sam.):

So what are the options?

OpenStudy (anonymous):